OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12 3y + 12 -3y + 12 3y - 12 -3y - 12

OpenStudy (a_clan):

-x - 3y = -12 multiply by -1 both sides x+3y=12 subtract 3y both sides x+3y-3y=-3y+12 or x = -3y + 12 (This is the required substitute)